# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ 3.4. Private Inference in WW-FL
In WW-FL, after the defined number of training iterations $T$ is completed, the MPC clusters serve as clusters for ML inference. Here, we again utilize PPML techniques to enable clients to privately query their clusters (Chaudhari et al., 2019, 2020; Knott et al., 2021; Mann et al., 2023). Consider the scenario where client $\mathsf{C}$ holding query ${Q}$ wants to use the inference service on a model $W$ that is secret-shared with a cluster $\mathcal{M}_{k}$. This is accomplished by $\mathsf{C}$ generating ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ using Share, followed by cluster servers in $\mathcal{M}_{k}$ running Predict on ${\langle W\rangle}_{\mathcal{M}_{k}}$ and ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ to generate the inference result in secret-shared form. Finally, $\mathcal{M}_{k}$ reveals the result to $\mathsf{C}$ using the Reveal protocol.
3.5. Abstraction of Existing FL Schemes
@@
The abstraction of WW-FL captures multiple existing FL frameworks, including standard FL with a single aggregator (McMahan et al., 2017), distributed aggregators (Fereidooni et al., 2021), and hierarchical FL schemes (Yang, 2021), as shown in Tab. 5. This abstraction not only simplifies comparisons but also facilitates advanced hybrid designs, including the integration of differential privacy.
@@
Table 5. Abstraction of existing FL schemes (cf. Tab. 1) in WW-FL (cf. Fig. 2). $\mathcal{S}$ – aggregation server(s), S.Agg. – secure aggregation, and (S.)Agg. – optional secure aggregation. See Tab. 2 for other notations.
@@
Standard FL with a single aggregator (Single $\mathcal{S}$(McMahan et al., 2017) is a variant of WW-FL, where each Layer III cluster $\mathcal{C}_{i}$ consists of only one client that also serves as the MPC cluster server $\mathcal{M}_{i}$ in Layer II. Thus, it is sufficient to conduct ML training without global model privacy concerns and then send the results to a single global server $\mathcal{G}$ in Layer I for aggregation. The case of distributed aggregators (Multi $\mathcal{S}$(Fereidooni et al., 2021) follows similarly, except that secure aggregation is performed at Layer I with multiple ($\mathsf{N}_{\mathcal{G}}>1$) global servers. Finally, existing hierarchical FL schemes (Yang, 2021) have a similar three-layer architecture as WW-FL, but use a single server at both the global and cluster level ($\mathsf{N}_{\mathcal{G}}=1$, $\mathsf{N}_{\mathcal{M}_{i}}=1$). While WW-FL employs PPML training at the cluster-server level, hierarchical FL uses secure aggregation. Additionally, clients in the hierarchical FL approach perform local model training, as opposed to only data sharing in WW-FL.
@@
Moreover, advanced hybrid designs in WW-FL could incorporate data from privileged clients, such as trusted consortium members, without hiding the global model. The client can train the model locally and share the gradients with the aggregation servers.
@@
Another hybrid could include a collaborative training model like SafeNet (Chaudhari et al., 2022), which resists poisoning attacks. SafeNet defines the global model as an ensemble of verified models, with the majority prediction determining the final output. This setup can replace an MPC cluster and its associated clients in WW-FL. The ensemble models can be aggregated or subjected to MPC-based clustering (Nguyen et al., 2022) to obtain a single model. The remaining process follows the WW-FL architecture, replacing operations in SafeNet on the plain model with corresponding MPC counterparts. Also, the training in the MPC clusters, which is resilient to poisoning attacks, can be categorized as attested training. The global servers can use a simple/mildly robust aggregation scheme to incorporate updates from these clusters, decreasing the likelihood of error propagation to higher layers.
4. Performance Evaluation
@@
We systematically and empirically evaluate the performance of WW-FL in terms of accuracy as well as computation and communication overhead. We use standard image classification tasks and build a prototype implementation using the MPC framework CrypTen (Knott et al., 2021). Future research can extend our evaluation to stronger security models (cf. §3.3), as well as more sophisticated ML models and training tasks.
4.1. Implementation

### v4

@@ 3.4 Private Inference in WW-FL
In WW-FL, after the defined number of training iterations $T$ is completed, the MPC clusters serve as clusters for ML inference. Here, we again utilize PPML techniques to enable clients to privately query their clusters [CRS20, KVH+21, MWCB23, NC23, HSW+25, SBBE25]. Consider the scenario where client $\mathsf{C}$ holding query Q{Q} wants to use the inference service on a model $W$ that is secret-shared with a cluster $\mathcal{M}_{k}$. This is accomplished by $\mathsf{C}$ generating ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ using the Share algorithm instantiated for cluster $\mathcal{M}_{k}$. Subsequently, the cluster servers in $\mathcal{M}_{k}$ running Predict on ${\langle W\rangle}_{\mathcal{M}_{k}}$ and ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ to generate the inference result in secret-shared form. Finally, $\mathcal{M}_{k}$ reveals the result to $\mathsf{C}$ using the Reveal protocol.
3.5 Abstraction of Existing FL Schemes
@@
The abstraction of WW-FL captures multiple existing FL frameworks, including standard FL with a single aggregator [MMR+17], distributed aggregators [TXLZ25], and hierarchical FL schemes [Yan21], as shown in Tab. 5. This abstraction not only simplifies comparisons but also facilitates advanced hybrid designs, including the integration of Differential Privacy [DP20, OA22].
@@
Table 5: Abstraction of existing FL schemes (cf. Tab. 1) in WW-FL (cf. Fig. 2). $\mathcal{S}$ – aggregation server(s), S.Agg. – secure aggregation, and (S.)Agg. – optional secure aggregation. See Tab. 2 for other notations.
@@
Standard FL with a single aggregator (Single $\mathcal{S}$[MMR+17] is a variant of WW-FL, where each Layer III cluster $\mathcal{C}_{i}$ consists of only one client that also serves as the MPC cluster server $\mathcal{M}_{i}$ in Layer II. Thus, it is sufficient to conduct ML training without global model privacy concerns and then send the results to a single global server $\mathcal{G}$ in Layer I for aggregation. The case of distributed aggregators (Multi $\mathcal{S}$[TXLZ25] follows similarly, except that secure aggregation is performed at Layer I with multiple ($\mathsf{N}_{\mathcal{G}}>1$) global servers. Finally, existing hierarchical FL schemes [Yan21] have a similar three-layer architecture as WW-FL, but use a single server at both the global and cluster level ($\mathsf{N}_{\mathcal{G}}=1$, $\mathsf{N}_{\mathcal{M}_{i}}=1$). While WW-FL employs PPML training at the cluster-server level, hierarchical FL uses secure aggregation. Additionally, clients in the hierarchical FL approach perform local model training, as opposed to only data sharing in WW-FL.
@@
Moreover, advanced hybrid designs in the WW-FL framework can accommodate privileged clients—such as trusted consortium members—who may contribute data without requiring the global model to remain hidden. In such cases, the client can train the model locally and secret-share the resulting gradients directly with the global aggregation servers $\mathcal{G}$.
@@
Another hybrid could involve collaborative training frameworks like SafeNet [CJO22], which offer robustness against poisoning attacks. SafeNet constructs the global model as an ensemble of individually verified models, using majority voting to determine the final prediction. In the context of WW-FL, this setup can serve as a substitute for a conventional MPC cluster and its associated clients. The ensemble models can then be either aggregated using standard techniques or clustered via MPC-based methods [NRC+22] to obtain a single representative model. From this point, the remaining operations conform to the WW-FL architecture, replacing SafeNet’s plaintext operations with their MPC counterparts. Notably, training performed within these MPC clusters—which inherently resist poisoning—can be classified as attested training. The global servers may then employ lightweight or mildly robust aggregation mechanisms to incorporate updates from these trusted clusters, thereby reducing the risk of error propagation to higher layers.
While our work primarily focuses on MPC for WW-FL components, we emphasize that our abstraction is general enough to accommodate other privacy-preserving techniques, including those based on homomorphic encryption (HE). For example, as discussed in §2, existing multi-party HE (MHE) schemes preserve global model privacy but face scalability limitations [FTP+21, SPT+21, XHX+23]. Within WW-FL, such an MHE-based setup can be treated as a single Layer II cluster. This allows the framework to leverage additional training data contributed by the MHE setup, instead of excluding it entirely. The main challenge in this integration lies in the Reshare algorithm, which would require converting between HE ciphertexts and MPC-style secret shares. Techniques proposed in works such as MP2ML [BCD+20] and Cheetah [HLHD22] can potentially be adapted to support these conversions, enabling seamless interoperability within WW-FL. We leave the integration of such HE-based clusters into WW-FL as an interesting direction for future work.
3.6 Example Configuration and Workflow in WW-FL
In this section, we outline the processing of the machine learning model within our WW-FL framework (cf. Fig. 2), using one specific configuration as an example from the various options listed in Tab. 4. This configuration, detailed in Tab. 6, involves three MPC clusters in Layer II, each using a different MPC protocol and global servers using the 2PC + helper protocol in CrypTen [KVH+21]. We use ${\langle x\rangle}_{\gamma}$ to denote the secret shares of the value $x$ held by all the parties in the set $\gamma$.
Table 6: Example configuration for WW-FL with three MPC Clusters in Layer II (cf. Fig. 2). ${\langle x\rangle}_{\gamma}$ represents the secret-shares of value $x$ held by all the parties in the set $\gamma$. In the 4PC setting [KPRS22], ${\lambda}_{x}^{j}$ denotes the input-independent shares, and ${\sf m}_{x}$ denotes the input-dependent share of $x$.
MPC Servers MPC Protocol Details
Layer I Global Servers: $\mathcal{G}$ 2PC + Helper [KVH+21] - semi-honest security. - $x=x_{1}+x_{2}$. - ${\langle x\rangle}_{\mathcal{G}}:\left(\mathsf{GS}_{1}:x_{1},\mathsf{GS}_{2}:x_{2}\right)$.
Layer II MPC Cluster 1: $\mathcal{M}_{1}$ 2PC Additive [DSZ15] - semi-honest security. - $x=x_{1}+x_{2}$. - ${\langle x\rangle}_{\mathcal{M}_{1}}:\left(\mathsf{CS}_{1}^{1}:x_{1},\mathsf{CS}_{1}^{2}:x_{2}\right)$.
Layer II MPC Cluster 2: $\mathcal{M}_{2}$ 3PC Replicated [AFL+16, MR18] - semi-honest security. - $x=x_{1}+x_{2}+x_{3}$. - ${\langle x\rangle}_{\mathcal{M}_{2}}$: $\mathsf{CS}_{2}^{1}:(x_{1},x_{2})$, $\mathsf{CS}_{2}^{2}:(x_{2},x_{3})$ $\mathsf{CS}_{2}^{3}:(x_{3},x_{1})$
Layer II MPC Cluster 3: $\mathcal{M}_{3}$ 4PC Replicated [KPRS22] - malicious security. - $x={\sf m}_{x}-{\lambda}_{x}^{1}-{\lambda}_{x}^{2}-{\lambda}_{x}^{3}$. - ${\langle x\rangle}_{\mathcal{M}_{3}}$: $\mathsf{CS}_{3}^{1}:({\sf m}_{x},{\lambda}_{x}^{1},{\lambda}_{x}^{2})$, $\mathsf{CS}_{3}^{2}:({\sf m}_{x},{\lambda}_{x}^{2},{\lambda}_{x}^{3})$, $\mathsf{CS}_{3}^{3}:({\sf m}_{x},{\lambda}_{x}^{3},{\lambda}_{x}^{1})$, $\mathsf{CS}_{3}^{4}:({\lambda}_{x}^{1},{\lambda}_{x}^{2},{\lambda}_{x}^{3})$.
The discussion follows the WW-FL training algorithm, detailed in Alg. 1 (FL-Train). For simplicity, we omit the process of clients secret-sharing their training datasets with the Layer II MPC clusters, as this can be done naively through the input-sharing protocol Share (see Tab. 3) of the underlying MPC protocol. Instead, we focus on the processing of the ML model among the MPC servers.
To enhance clarity in our discussion, we have divided the FL-Train algorithm into five steps, presented in the order of execution. We describe the workflow throughout these steps, and for reference, we have included the corresponding line numbers from Alg. 1 in the format ${\sf L}123$. For the remainder of this section, we assume that all operations are performed over a finite $\ell$-bit ring, as required by the underlying MPC protocols. Additionally, to simplify the presentation, we omit the discussion of communication optimizations involving shared-key setups used in these protocols, and we refer the reader to [KVH+21, DSZ15, AFL+16, MR18, KPRS22] for further details.
[…]
In this step, the initial global model $W_{0}$, likely owned by an organization or corporate entity, is secret-shared among the two global servers $\mathsf{GS}_{1},\mathsf{GS}_{2}\in\mathcal{G}$. The model owner uses CrypTen’s [KVH+21] input sharing protocol Share to secret-share $W_{0}$, resulting in ${\langle W_{0}\rangle}_{\mathcal{G}}$, i.e., $W_{0}={\langle W_{0}\rangle}_{\mathsf{GS}_{1}}+{\langle W_{0}\rangle}_{\mathsf{GS}_{2}}$.
[…]
This step involves resharing the global model $W_{t-1}$ for the current iteration $t$, to the three MPC clusters. At a high level, each global server $\mathsf{GS}_{i}$ acts as an input party for the MPC protocol within cluster $\mathcal{M}_{j}$ and secret-shares its share of the global model. Specifically, for global model $x=W_{t-1}$, $\mathsf{GS}_{i}$ secret-shares $x_{i}={\langle W_{t-1}\rangle}_{\mathsf{GS}_{i}}$, among the cluster servers $\mathcal{M}_{j}$, represented as ${\langle x_{i}\rangle}_{\mathcal{M}_{j}}$.
Due to the linearity of MPC schemes, the cluster servers can locally add these shares. This results in a secret sharing of $x=W_{t-1}$ across the servers in cluster $\mathcal{M}_{j}$, represented as ${\langle x\rangle}_{\mathcal{M}_{j}}={\langle x_{1}\rangle}_{\mathcal{M}_{j}}+{\langle x_{2}\rangle}_{\mathcal{M}_{j}}$.
MPC Cluster $\mathcal{M}_{1}$
This cluster uses 2-out-of-2 additive secret sharing with semi-honest security [DSZ15]. To generate ${\langle x\rangle}_{\mathcal{M}_{1}}$ with $x=W_{t-1}$, the servers perform the following steps:
  • Each $\mathsf{GS}_{i}\in\mathcal{G}$ samples random shares $x_{i,1}$ and $x_{i,2}$ such that $x_{i}=x_{i,1}+x_{i,2}$.
  • Each $\mathsf{GS}_{i}$ sends $x_{i,j}$ to cluster server $\mathsf{CS}_{1}^{j}\in\mathcal{M}_{1}$.
  • Each $\mathsf{CS}_{1}^{j}$ locally computes ${\langle x\rangle}_{\mathsf{CS}_{1}^{j}}=x_{1,j}+x_{2,j}$.
MPC Cluster $\mathcal{M}_{2}$
This cluster is instantiated with a three-party replicated secret sharing with semi-honest security [AFL+16, MR18]. To generate ${\langle x\rangle}_{\mathcal{M}_{2}}$ with $x=W_{t-1}$, the servers proceed as follows:
  • Each $\mathsf{GS}_{i}\in\mathcal{G}$ samples random shares $x_{i,1}$, $x_{i,2}$ and $x_{i,3}$ such that $x_{i}=x_{i,1}+x_{i,2}+x_{i,3}$.
  • $\mathsf{GS}_{i}$ sends the pair $(x_{i,j},\,x_{i,j\%3+1})$ to $\mathsf{CS}_{2}^{j}\in\mathcal{M}_{2}$ as its share in ${\langle x_{i}\rangle}_{\mathcal{M}_{2}}$.
  • Servers in $\mathcal{M}_{2}$ locally compute ${\langle x\rangle}_{\mathcal{M}_{2}}={\langle x_{1}\rangle}_{\mathcal{M}_{2}}+{\langle x_{2}\rangle}_{\mathcal{M}_{2}}$. Specifically, each $\mathsf{CS}_{2}^{j}$ computes its share of ${\langle x\rangle}_{\mathcal{M}_{2}}$ as
    ${\langle x\rangle}_{\mathsf{CS}_{2}^{j}}=(x_{1,j}+x_{2,j},x_{1,j\%3+1}+x_{2,j\%3+1}).$
MPC Cluster $\mathcal{M}_{3}$
This cluster consists of four servers and utilizes the four-party (4PC) replicated secret sharing with malicious security from Tetrad [KPRS22]. Note that Tetrad uses function-dependent preprocessing to achieve faster online phase. Consequently, the cluster servers in $\mathcal{M}_{3}$ generate all the input-independent $\lambda$} values during the preprocessing phase. Our discussion begins at this point, assuming that the preprocessing by the cluster servers has been completed. From here, the servers proceed as follows:
  • Cluster servers send the following $\lambda$} shares to $\mathsf{GS}_{i}\in\mathcal{G}$:
    $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{3},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x_{i}}^{1}$
    $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{2},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x_{i}}^{2}$
    $\mathsf{CS}_{3}^{2},\mathsf{CS}_{3}^{3},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x_{i}}^{3}$
  • $\mathsf{GS}_{i}$ accepts the $\lambda$} values that form the majority and computes ${\lambda}_{x_{i}}={\lambda}_{x_{i}}^{1}+{\lambda}_{x_{i}}^{2}+{\lambda}_{x_{i}}^{3}$.
  • $\mathsf{GS}_{i}$ sends ${\sf m}_{x_{i}}=x_{i}-{\lambda}_{x_{i}}$ to $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{2}$, and $\mathsf{CS}_{3}^{3}$, thus completing the input sharing to generate ${\langle x_{i}\rangle}_{\mathcal{M}_{3}}$.
  • Servers in $\mathcal{M}_{3}$ locally compute ${\langle x\rangle}_{\mathcal{M}_{3}}={\langle x_{1}\rangle}_{\mathcal{M}_{3}}+{\langle x_{2}\rangle}_{\mathcal{M}_{3}}$.
[…]
This step involves MPC-based ML training at the cluster servers, using secret shares of the current global model, represented as ${\langle W_{t-1}\rangle}_{\mathcal{M}_{j}}$, along with the data from the clients associated with the cluster in the current iteration, denoted by ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}$. The servers then execute the underlying PPML training algorithm, Train, and obtain secret shares of the updated model, denoted by ${\langle W_{t}\rangle}_{\mathcal{M}_{j}}$.
3.6.4 Step IV ($1{\sf L}\ref${line:fltrain-step-d}): Resharing of Local Model
Once each cluster receives its updated local model for iteration $t$ in secret-shared form ${\langle W_{t}\rangle}_{\mathcal{M}_{j}}$, the next step is to reshare this model among the global servers $\mathcal{G}$. This process is similar to Step II above, with the key difference being that the roles of the source and destination of the secret sharing are now reversed. We detail the steps for each of the three Layer II clusters next.
MPC Cluster $\mathcal{M}_{1}$
To generate ${\langle x\rangle}_{\mathcal{G}}$ with $x=W_{t}^{1}$, the cluster servers in $\mathcal{M}_{1}$ perform the following steps:
  • Each $\mathsf{CS}_{1}^{i}\in\mathcal{M}_{1}$ samples random shares $x_{i,1}$ and $x_{i,2}$ such that $x_{i}=x_{i,1}+x_{i,2}$.
  • Each $\mathsf{CS}_{1}^{i}$ sends $x_{i,j}$ to global server $\mathsf{GS}_{j}\in\mathcal{G}$.
  • Each $\mathsf{GS}_{j}$ locally computes ${\langle x\rangle}_{\mathsf{GS}_{j}}=x_{1,j}+x_{2,j}$.
MPC Cluster $\mathcal{M}_{2}$
The resharing process in cluster $\mathcal{M}_{2}$, which uses a three-party replicated sharing scheme where $x=W_{t}^{2}$ is represented as $x=x_{1}+x_{2}+x_{3}$, closely resembles that of cluster $\mathcal{M}_{1}$ with a few minor differences.
First, cluster $\mathcal{M}_{2}$ involves three servers, $\mathsf{CS}_{2}^{i}$, compared to the two servers in $\mathcal{M}_{1}$. Consequently, each global server $\mathsf{GS}_{j}$ will receive secret shares of three values instead of two. Second, the local computation for a secret-shared value $x$ in $\mathcal{M}_{2}$ involves adding three shares: ${\langle x\rangle}_{\mathsf{GS}_{j}}=x_{1,j}+x_{2,j}+x_{3,j}$, rather than just two as in $\mathcal{M}_{1}$.
Additionally, because $\mathcal{M}_{2}$ operates under semi-honest security, an optimization can be applied. Specifically, $\mathsf{CS}_{2}^{1}$ can directly secret-share the sum $(x_{1}+x_{2})$, while $\mathsf{CS}_{2}^{2}$ secret-shares $x_{3}$, effectively eliminating one instance of secret-sharing.
MPC Cluster $\mathcal{M}_{3}$
Since this cluster uses a maliciously secure four-party protocol from Tetrad [KPRS22], we cannot directly follow the approach where each cluster server $\mathsf{CS}_{3}^{i}$ secret-shares its share of $x=W_{t}^{3}$ with the global servers in $\mathcal{G}$. However, because the protocol uses replicated secret sharing, each share is held by three of the four parties, with at most one being corrupt. Therefore, instead of a single cluster server sharing its own secret, the share can be “jointly” shared by three cluster servers. This ensures that a majority of correct values is always available at the global server $\mathsf{GS}_{j}$. The remaining steps are similar to those for other clusters and are described below.
  • Cluster servers “jointly” secret share the following shares among the two global servers in $\mathcal{G}$:
    $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{3},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x}^{1}$
    $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{2},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x}^{2}$
    $\mathsf{CS}_{3}^{2},\mathsf{CS}_{3}^{3},\mathsf{CS}_{3}^{4}$ $:{\lambda}_{x}^{3}$
    $\mathsf{CS}_{3}^{1},\mathsf{CS}_{3}^{2},\mathsf{CS}_{3}^{3}$ $:{\sf m}_{x}$
  • For each of the four instances, $\mathsf{GS}_{i}$ accepts the share that forms the majority.
  • Servers in $\mathcal{G}$ locally compute ${\langle x\rangle}_{\mathcal{G}}={\langle{\sf m}_{x}\rangle}_{\mathcal{G}}-{\langle{\lambda}_{x}^{1}\rangle}_{\mathcal{G}}-{\langle{\lambda}_{x}^{2}\rangle}_{\mathcal{G}}-{\langle{\lambda}_{x}^{3}\rangle}_{\mathcal{G}}$.
[…]
In this example, the global servers are implemented using CrypTen [KVH+21], and the secure aggregation can be executed using the WW-FL code, which provides an implementation of secure aggregation in CrypTen. However, the global servers are not limited to CrypTen. They can also be instantiated with other MPC protocols. For example, SafeFL [GMS+23] offers a PyTorch connector plugin for using PyTorch with the MP-SPDZ [Kel20] framework, which supports a range of efficient MPC protocols.
Similarly, while we use three distinct MPC protocols with a small number of parties to instantiate the MPC clusters in Layer II, our framework is not limited to these specific protocols. The technique of resharing secret shares is well-known in the MPC literature and has been used in several works such as [AAG+25, BCD+20], though it is less common in the context of federated learning. For example, the “converter unit” in FANNG-MPC [AAG+25] introduces methods to convert secret shares from one set of $n$ parties in a dishonest majority setting with malicious security to another set of $m\neq n$ parties in the same corruption setting. Similarly, MP2ML [BCD+20] presents a scheme for converting between HE (using the CKKS scheme) and MPC to enable private ML inference. These works highlight the flexibility of our WW-FL framework, which enables global federated learning and integrates various existing techniques.
4 Performance Evaluation
@@
We systematically and empirically evaluate the performance of WW-FL in terms of accuracy as well as computation and communication overhead. We use standard image classification tasks and build a prototype implementation [^6] https://encrypto.de/code/WW-FL using the MPC framework CrypTen [KVH+21]. Future research can extend our evaluation to stronger security models (cf. §3.3), as well as more sophisticated ML models and training tasks.
4.1 Implementation